# Reduction-induced the Variation of Partial Von Neumann Entropy

### v1

@@ Abstract
Whether from theoretical research or from practical research and applications, bipartite mixed-state quantum entanglement (QE) and its measures hold significant theoretical and practical importance. Widely recognized measures such as Entanglement of Formation and Relative Entropy of Entanglement, involve quantifying the minimum QE and reflect the inherently complex nature of their computation. In contrast, Concurrence is applicable only to $2\times 2$ quantum systems, while Negativity is limited to $2\times 2$ and $2\times 3$ quantum systems. In this context, we propose a novel method termed Reduction-induced the Variation of Partial Von Neumann Entropy (RIVPVNE) to quantify QE in bipartite mixed states. The computational complexity associated with this method is minimal. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, combined with straightforward computation, facilitates the exploration of its specific potential applications. Subsequently, we employ this method to investigate the organization and structure of some well-known QEs, analyzing its characteristics through both its physical implication and mathematical structure in comparison to the existing measures for bipartite QE measures, such as Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation and Relative Entropy of Entanglement.
Keywords:Quantum Entanglement (QE); Bipartite Mixed-state QE; Reduction-induced the Variation of Partial Von Neumann Entropy (RIVPVNE)
@@ I Introduction
In the rapidly developing field of quantum information, quantum entanglement (QE) has become one of the most mysterious and key concepts, triggering a wide range of research interests. The concept of QE was first proposed and described as a strange phenomenon of quantum mechanics that called into question the completeness of the quantum theory by the famous physicists Einstein, Podolsky, and Rosen (EPR) epr and Schrödinger sd . Later on, Bell realized that QE can produce experimentally verifiable deviations from classical physics in quantum mechanics bell . However, even to this day, its mystery still challenges our understanding of the laws of nature.
@@
The essence of QE is that when two or more quantum systems interact with each other, their states cannot be described individually but only as a whole. This ”non-local” characteristic has led to revolutionary applications such as quantum cryptography kk , quantum teleportation ch , quantum superdense coding sj , and measurement-based quantum computation rr . Therefore, a deeper understanding and accurate metric of QE have become important prerequisites for the development of quantum information science and quantum technology.
@@
Nonseparable bipartite systems exhibit QE epr ; sd , which can be quantified using various measures. These include the Partial Von Neumann Entropy ch1 ; ch2 , applicable only to bipartite pure states, and Concurrence, which is confined to $2\times 2$ quantum systems wk ; au ; ka ; mf ; xjw , and Negativity, which is applicable only to $2\times 2$ and $2\times 3$ quantum systems K.Z ; gv . Additionally, Entanglement of Formation ch1 ; ch2 and Relative Entropy of Entanglement vv1 ; vv2 for bipartite quantum states, involve quantifying the minimum QE, and their assessments generally entail considerable computational burdens, especially for bipartite mixed-state QE.
@@
Mixed states are more reflective of real-world scenarios, as factors such as environmental conditions, quantum measurements, and thermal influences can lead quantum systems to exist in mixed states. Consequently, the measures of QEs in mixed states are often more directly applicable to complex systems, aligning better with experimental conditions and practical applications. In general, QEs in pure states are more intuitive and easier to comprehend, while those in mixed states are more closely related to practical applications. In practical research and applications, bipartite mixed-state QE and its measures hold significant practical importance, besides their theoretical significance. The organization and structure of mixed-state QE is more complicated and less well understood than that of pure-state QE. Quantifying QEs in bipartite mixed states is more complex and challenging due to the contributions from classical probability distributions, making the definition and computation of QEs more intricate.
@@
In this context, our research aims to explore new metrics for bipartite mixed-state QE. We propose a novel method termed Reduction-induced the Variation of Partial Von Neumann Entropy to quantify QE in bipartite mixed states. The computational complexity associated with this method is considerably low. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, combined with straightforward computation, facilitates the exploration of its specific potential applications. Subsequently, we employ this method to investigate the organization and structure of some well-known QEs, analyzing its characteristics through both its physical implication and mathematical structure in comparison to the existing bipartite QE measures, such as Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation and Relative Entropy of Entanglement.
@@ II.3 Relative Entropy of Entanglement for Bipartite QE
$\operatorname{REE}=S(\rho|\sigma)=\min_{\sigma\in S}[-\operatorname{Tr}\rho\log_{2}{\sigma}-S(\rho)],$ (8)
where S represents a set of separable bipartite density matrix. On one hand, Relative Entropy of Entanglement measures the shortest ”distance” from a given quantum entangled state to the set of separable states, which is the geometric interpretation of this method and reflects that the involving computation is inherently complex. On the other hand, while for bipartite pure QE states, Relative Entropy of Entanglement (REE) simplifies to the Partial Von Neumann Entropy (PVNE)
$\operatorname{REE}$ $=S(\rho|\sigma)=\min_{\sigma\in S}[-\operatorname{Tr}\rho\log_{2}{\sigma}]$ (9)
@@ II.4 Negativity for $2\times 2$ and $2\times 3$ quantum systems
Summary: Due to the locality of the aforementioned five kinds of bipartite QE measures, symmetry about different partitions is generally observed only for symmetric states. For mixed-state QE, the widely recognized measures for quantifying QE in bipartite mixed states, such as Entanglement of Formation (EOF) and Relative Entropy of Entanglement (REE), involve determining the minimum QE required for a given state, reflecting the inherent complexity of their computation. In contrast, Concurrence is applicable only to $2\times 2$ quantum systems, while Negativity is limited to $2\times 2$ and $2\times 3$ quantum systems. The organization and structure of mixed-state QE is more complicated and less well understood than that of pure-state QE. Quantifying QE in mixed states poses additional challenges due to the contributions from classical probability distributions, which complicate both the definition and computation of QE measures. In this context, our research aims to explore new measures for bipartite mixed-state QE in detail, as outlined below.
III Reduction-induced the Variation of Partial Von Neumann Entropy (RIVPVNE)
In any bipartite quantum system, all the information is contained within its wave function or density matrix. When a subsystem is obtained by partial tracing or averaging out the other subsystem freedom, some information is inevitably lost. Therefore, the concept of Reduction-induced the Variation of Partial Von Neumann Entropy is presented to quantify bipartite QE, which holds certain importance and practical significance. To elaborate further, let’s consider a bipartite quantum system. In the bipartite density matrix $\rho$, the density matrix of the $i$-th $(i=1,2)$ party, denoted as the bipartite $i$-th party density matrix, only preserves the information of the $i$-th party from $\rho$. This particular density matrix is represented as $\rho(i)$ different from its reduced density matrix $\rho_{i}$. Reduction-induced the Variation of Von Neumann Entropy (RIVPVNE) of one-sided density matrix is defined as follows:
@@ III Reduction-induced the Variation of Partial Von Neumann Entropy (RIVPVNE)
which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to bipartite states. Notably, the computational complexity associated with this method is minimal. Moreover, it has a wide range of applicability. The intuitive and clear physical picture, combined with straightforward computation, facilitates the exploration of its specific potential applications. For bipartite pure-state QE, this measure reduces to Partial Von Neumann Entropy
@@
Due to the locality of this measure, in general only for symmetric states, it has symmetry about different partition. Subsequently, we employ it to investigate the organization and structure of some well-known QEs, analyzing its characteristics through both its physical implication and mathematical structure in comparison to existing bipartite QE measures, such as Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation and Relative Entropy of Entanglement, as detailed below.
IV Its Applications and Relative Characteristics in comparison to existing bipartite QE measures
@@
Consequently, the Reduction-induced the Variation of Partial Von Neumann Entropy, as well as the Entanglement of Formation and Relative Entropy of Entanglement, simplifies to the Partial Von Neumann Entropy:
@@
This highlights the interference properties and indicates a clear and intuitive relationship between QE and quantum coherence. The previously mentioned measures are illustrated in Fig.1. For the specified pure states, it is observed that when $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, all the measures satisfy: $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and for the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$ hold. Clearly, for the specified pure two-qubit systems, the monotonicities of these measures are fully consistent. The Reduction-induced Variation of Partial Von Neumann Entropy, the Entanglement of Formation, and the Relative Entropy of Entanglement reduce to the Partial Von Neumann Entropy, which is slightly different from Concurrence and Negativity, stemming from their different mathematical structure and physical implication.
Fig. 1: Reduction-induced the Variation of Partial Von Neumann Entropy as well as other QE measures such as the Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for the values $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and in the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$.
@@
The first term corresponds to a QE state with a probability of $\sin^{2}\alpha$, while the second term represents a non-entangled state with a probability of $\cos^{2}\alpha$. Consequently, the bipartite QE cannot exceed $\sin^{2}\alpha$. The corresponding single-site reduced density matrix and bipartite single-site density matrix are
@@
Next, the Reduction-induced the Variation of Partial Von Neumann Entropy (RIVPVNE) is computed as
@@
which represents the upper limit of QE as derived from the mathematical structure of the system’s density matrix, and the Negativity expressed as
@@
The five aforementioned measures are illustrated in Fig.2. Notably, when $\alpha=\frac{\pi}{2}$, the following relations hold: $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. For values $\alpha=0.8$, and $2.341$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ hold; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, we find $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, the relations become $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$. Clearly, for the specified mixed two-qubit quantum systems, the monotonic behavior of these five measures is consistent. Specifically, the differences in the values of Reduction-induced the Variation of Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence, and Negativity arise from their distinct mathematical structures and physical implications. In terms of the bipartite QE degree metric for the specified mixed states, it is concluded that Reduction-induced the Variation of Partial Von Neumann Entropy, Entanglement of Formation, and Negativity are preferable to Relative Entropy of Entanglement and Concurrence. In terms of bipartite QE degree metric of the specified mixed states, Reduction-induced the Variation of Partial Von Neumann Entropy, Entanglement of Formation, and Negativity are superior to Relative Entropy of Entanglement and Concurrence.
@@
Fig. 2: Reduction-induced the Variation of Partial Von Neumann entropy as well as other QE measures such as Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$. For values $\alpha=0.8$, and $2.341$, $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ ; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$.

### v2

# Reduction-induced Variation of Partial Von Neumann Entropy

@@ Abstract
The organization and structure of bipartite mixed-state quantum entanglement (QE) is more complicated and less well understood than that of bipartite pure-state QE. Bipartite mixed-state QEs and their measures play a crucial role in both theory and practical applications. Existing measures comprise two classes: one class involves quantifying the minimum QE and reflect the inherently complex nature of their computation, while the other is only applicable to highly limited-dimensional quantum systems. In this context, we propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) to quantify QE in any bipartite states especially for any bipartite mixed-states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states. This method exhibits minimal computational complexity and broad applicability. Its intuitive and clear physical representation, combined with easy computation and wide applicability, facilitates exploring its potential applications. Furthermore, we present examples to demonstrate the superiorities of this method in identifying bipartite QE by comparing with other existing bipartite mixed-state QE measures through both their physical implications and mathematical structures.
Keywords:Quantum Entanglement (QE); Bipartite Mixed-state QE; Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
@@ I Introduction
In 1935, “Spooky Action at a Distance”described as a strange phenomenon of quantum mechanics calls into question the completeness of the quantum theory by the famous physicists Einstein, Podolsky and Rosen (EPR) epr . And then in the same year, the concept of quantum entanglement (QE) was first proposed clearly by Schrödinger sd , through rigorous mathematical formalism, he demonstrated that particles in such entangled states cannot be individually described but must be treated as an integral composite system. Subsequently, in 1964, Bell realized that QE can produce experimentally verifiable deviations from classical physics in quantum mechanics bell . However, even to this day, its mystery still challenges our understanding of the laws of nature.
@@
The type or degree of QEs are often directly related to their revolutionary applications including quantum cryptography kk ; ng , quantum teleportation ch ; rc ; nr , quantum superdense coding nr ; sj , measurement-based quantum computation rr ; fa and quantum enhanced sensing tc ; rcp ; lp . Therefore, a deeper understanding and accurate metric of QE have become important prerequisites for the development of quantum information science and quantum technology.
@@
Nonseparable bipartite systems exhibit QE epr ; sd , which can be quantified using various measures. These include the Partial Von Neumann Entropy ch1 ; ch2 , applicable only to bipartite pure states, and Concurrence, which is confined to $2\times 2$ quantum systems wk ; au ; ka ; mf ; xjw , and Negativity, which is applicable only to $2\times 2$ and $2\times 3$ quantum systems K.Z ; gv . Additionally, Entanglement of Formation ch1 ; ch2 and Relative Entropy of Entanglement vv1 ; vv2 for bipartite quantum states, involve quantifying the minimum QE, and their assessments generally entail considerable computational burdens, especially for bipartite mixed-state QE. In general, QEs in pure states are more intuitive and easier to comprehend, while those in mixed states are more complicated and less well understood than those of pure-state QEs. Quantifying QEs in bipartite mixed states is more complex and challenging due to the contributions from classical probability distributions, making the definition and computation of QEs more intricate.
@@
Mixed states are more reflective of real-world scenarios, as factors such as environmental conditions, quantum measurements, and thermal influences can lead quantum systems to exist in mixed states. Consequently, the measures of QEs in mixed states are often more directly applicable to complex systems, aligning better with experimental conditions and practical applications. Hence, in practical research and applications, bipartite mixed-state QE and its measures hold significant practical importance, besides the theoretical significance of their underlying structure of quantum mechanics.
@@
In this context, our research aims to explore new metrics for bipartite mixed-state QE. We propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states. The computational complexity associated with this method is considerably low, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. Subsequently, we provide examples to demonstrate the superiorities of this method in bipartite QE identification through a comparison with other established bipartite mixed-state QE measures including Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation, and Relative Entropy of Entanglement, considering both their physical implications and mathematical structures.
@@ II.3 Relative Entropy of Entanglement for Bipartite QE
$\operatorname{REE}=S(\rho|\sigma)=\min_{\sigma\in\mathcal{S}}[-\operatorname{Tr}\rho\log_{2}{\sigma}-S(\rho)],$ (8)
where $\mathcal{S}$ represents a set of separable bipartite density matrix. On one hand, Relative Entropy of Entanglement measures the shortest ”distance” from a given quantum entangled state to the set of separable states, which is the geometric interpretation of this method and reflects that the involving computation is inherently complex. On the other hand, while for bipartite pure QE states, Relative Entropy of Entanglement (REE) simplifies to the Partial Von Neumann Entropy (PVNE)
$\operatorname{REE}$ $=S(\rho|\sigma)=\min_{\sigma\in\mathcal{S}}[-\operatorname{Tr}\rho\log_{2}{\sigma}]$ (9)
@@ II.4 Negativity for $2\times 2$ and $2\times 3$ quantum systems
Summary: Due to the locality of the aforementioned five kinds of bipartite QE measures, symmetry about different partitions is generally observed only for symmetric states. The widely recognized measures for quantifying bipartite mixed-state QE, such as Entanglement of Formation and Relative Entropy of Entanglement, involve determining the minimum QE required for a given state, reflecting the inherent complexity of their computation. While Concurrence is applicable only to $2\times 2$ quantum systems and Negativity limited to $2\times 2$ and $2\times 3$ quantum systems. The organization and structure of bipartite mixed-state QE is more complicated and less well understood than that of pure-state QE. Quantifying QE in mixed states poses additional challenges due to the contributions from classical probability distributions, which complicate both the definition and computation of QE measures. In this context, our research aims to explore new measures for bipartite mixed-state QE in detail, as outlined below.
III Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
In any bipartite quantum system, all the information is contained within its wave function or density matrix. When a subsystem is obtained by partial tracing or averaging out the other subsystem freedom, some information is inevitably lost. Therefore, the concept of Reduction-induced Variation of Partial Von Neumann Entropy is presented to quantify bipartite QE, which holds certain importance and practical significance. To elaborate further, let’s consider a bipartite quantum system. In the bipartite density matrix $\rho$, the density matrix of the $i$-th $(i=1,2)$ party, denoted as the bipartite $i$-th party density matrix, only preserves the information of the $i$-th party from $\rho$. This particular density matrix is represented as $\rho(i)$ different from its reduced density matrix $\rho_{i}$. Reduction-induced Variation of Von Neumann Entropy (RIVPVNE) of one-sided density matrix is defined as follows:
@@ III Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states especially for any bipartite mixed states. This measure vanishes if and only if the state is separable, and on average, does not experience an increase under stochastic Local Operations and Classical Communications, thereby establishing itself as an entanglement monotone. Notably, the computational complexity associated with this method is minimal, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical picture, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. For any bipartite pure-state QE, this measure reduces to Partial Von Neumann Entropy
@@
Due to the locality of this measure, in general only for symmetric states, it has symmetry about different partition. Subsequently, we provide examples to demonstrate the superiorities of this method in identifying QE through a comparison with other existing bipartite QE measures, such as Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation and Relative Entropy of Entanglement, considering both their physical implications and mathematical structures, as detailed below.
IV Its Superiorities in comparison to existing bipartite QE measures
@@
Consequently, the Reduction-induced Variation of Partial Von Neumann Entropy, as well as the Entanglement of Formation and Relative Entropy of Entanglement, simplifies to the Partial Von Neumann Entropy:
@@
This highlights the interference properties and indicates a clear and intuitive relationship between QE and quantum coherence. The previously mentioned measures are illustrated in Fig.1. For the specified pure states, it is observed that when $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, all the measures satisfy: $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and for the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$ hold. Clearly, for the specified pure two-qubit systems, the monotonicities of these measures are fully consistent. The Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation and Relative Entropy of Entanglement reduce to the Partial Von Neumann Entropy, but exhibit slight differences from Concurrence and Negativity due to variations in their physical implications and mathematical structures.
Fig. 1: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as the Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for the values $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and in the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$.
@@
The first term corresponds to a QE state with a probability of $\sin^{2}\alpha$, while the second term represents a non-entangled state with a probability of $\cos^{2}\alpha$. Consequently, the bipartite QE is less than $\sin^{2}\alpha$. The corresponding single-site reduced density matrix and bipartite single-site density matrix are
@@
Next, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is computed as
@@
which represents the upper limit of QE as derived from the mathematical structure of the system’s density matrix and reflect that they may not accurately reflect the degree of mixed-state QE, and the Negativity expressed as
@@
The five aforementioned measures are illustrated in Fig.2. Notably, when $\alpha=\frac{\pi}{2}$, the following relations hold: $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. For values $\alpha=0.8,2.341$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ hold; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, we find $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, the relations become $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$. Clearly, for the specified mixed two-qubit quantum systems, the monotonic behavior of these five measures is consistent. Specifically, the differences in the values of Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence, and Negativity arise from their distinct physical implications and mathematical structures. Regarding the bipartite QE degree metric for the specified mixed states, it is concluded that the Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation, and Negativity exhibit advantages over the Relative Entropy of Entanglement and Concurrence. When considering bipartite QE degree metric in mixed states, the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy, is superior to Entanglement of Formation, Negativity, Relative Entropy of Entanglement and Concurrence.
@@
Fig. 2: Reduction-induced Variation of Partial Von Neumann entropy as well as other QE measures such as Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$. For values $\alpha=0.8,2.341$, $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ ; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$.